# Understanding and Mitigating Language Confusion in LLMs

### v1

@@ 1 Introduction
Large language models (LLMs) are increasingly used in a variety of applications across the globe Kaddour et al. (2023). While early LLMs focused on English Joshi et al. (2020); Hu et al. (2020), recent models are more multilingual. Nevertheless, LLMs do not provide equal utility to speakers of non-English languages due to higher latency, increased costs, and reduced performance Ahia et al. (2023); Asai et al. (2023); Held et al. (2023).
@@
Take an Arabic prompt as an example: an LLM may inappropriately respond fully in English (full-response confusion), produce some lines in the desired language and some in another language (line-level confusion), or sporadically insert single words or phrases in another language (word-level confusion). Figure 1 shows example errors.
We investigate language confusion on the line and word level [^3] For word-level confusion, we focus on English words given the English-centricity of most LLMs and the difficulty of reliably reproducing and measuring non-English word switches. Albeit rarely, occurrences of non-English word-level confusions do happen, as shown in Figure A1. in two practical settings: a) Monolingual generation, where a user queries the LLM in a given language, implicitly requesting an answer in the same language; and b) cross-lingual generation, where a user explicitly instructs a model to generate text in a different language.
@@ 2.2 Language Confusion Metrics
Off-the-shelf tools do not support word-level LID and LLMs only achieve 79–86 F1 detecting word-level code-switching Zhang et al. (2023), too low for use as automatic evaluators. Observing that word-level confusion in top LLMs mainly occurs with English, we opt for a heuristic evaluation restricted to a subset of languages where it achieves high precision. We thus focus on non-Latin script languages, where Latin script words are trivially detected. To avoid natural code-switching false positives, [^5] E.g., Japanese users may use English acronyms: {CJK}UTF8minAI に関する記事を書きます。(“Write an article about AI.”) we check for English words that do not typically occur in target language text. [^6] We source words from https://gist.github.com/WChargin/8927565, which is based on the Linux dictionary usually stored in /usr/share/dict/words. To reduce false positives, we exclude capitalized words from the dictionary (which are often proper nouns or acronyms). We evaluate word-level confusion in Arabic (ar), Hindi (hi), Japanese (ja), Korean (ko), Russian (ru), and Simplified Chinese (zh).
@@
A response is only correct when entirely in the correct language, as even one instance of language confusion can damage intelligibility and cause a jarring user experience. We calculate binary metrics to indicate whether a response contains any instance of a) a line in an incorrect language and b) an isolated English word/phrase for languages using non-Latin scripts. These main metrics in our evaluation are defined below.
@@
where $R$ is the set of all responses and $E_{L}$ the set of responses that contain line-level errors. [^7] For completeness, we show results for “line-level language accuracy” (the fraction of lines in the correct language across all responses of an LLM) in Tables A1 and A2 at §A.1
Word-level pass rate (WPR): percentage of responses where all words are in the desired language. We exclude responses with line-level errors as most line-level errors would also be counted as word-level errors, making it difficult to disentangle the two error types. Detecting word-level language confusion is highly challenging so we only evaluate on target languages that use a non-Latin script and look for erroneous English words—the most common type of word-level language confusion.
@@ 2.4 Data Filtering and Processing
As LID tools underperform on short sequences and non-standard text, we manually filter out: a) examples answerable with a single word/phrase; b) multiple-choice questions and prompts asking for lists; c) prompts requiring code generation, math equations, or data formats such as HTML. For datasets where completions are available, we prioritize the prompts with the longest completions (i.e. more than 4 words).
@@
Table 4: Average word-level pass rate (WPR). For a detailed by-language breakdown, see Table A3.
@@ 3 Experiments
We evaluate the following LLMs covering various scales and model families: Llama 2 70B Instruct Touvron et al. (2023), Llama 3 70B Instruct, [^12] https://ai.meta.com/blog/meta-llama-3/ Command R (35B parameters), [^13] https://cohere.com/blog/command-r Command R+ (104B parameters), [^14] https://cohere.com/blog/command-r-plus-microsoft-azure Mixtral 8x7B Jiang et al. (2024), Mistral Large, [^15] https://mistral.ai/news/mistral-large/ GPT-3.5 Turbo (gpt-3.5-turbo-012524; Brown et al., 2020) and GPT-4 Turbo (gpt-4-turbo-040924; Achiam et al., 2023). For Llama and Command models, we also evaluate base versions (see §4.5). We generate at most 100 tokens per prompt using nucleus sampling with $p=0.75$ and $T=0.3$.
LPR, WPR, and LCPR results are in Tables 3, 4, and A4, respectively. [^16] We find results show little variance across runs (see §A.4). We show language confusion examples for different models in Table A5.
@@
Command and GPT models perform well on average on the line level (LPR in $[98.6,99.3]$), but Llama and Mistral models struggle to consistently generate text in the correct language (LPR in $[48.3,73.0]$). Mistral Large is better on some European languages while Llama models exhibit language confusion even for high-resource languages such as German. Most models have WPR within the same range, though Mixtral 8x7B is considerably worse. GPT-4 Turbo is strongest on LCPR, with GPT-3.5 Turbo and Command R+ tied for second. Llama models’ preference for English responses leads to very low LCPR.
@@
In the challenging cross-lingual setting, the best models have LPRs in the low 90s. Command and OpenAI models are best, with Command R+ best overall. Llama models perform poorly: both scoring in the 30s due to a tendency to respond in English. Mistral Large is worse than Mixtral, even in European languages. For non-Latin script languages, Command R+ and OpenAI models are again best. LCPR shows a similar trend as in the monolingual task, albeit with lower scores: GPT-3.5 Turbo performs best, with Command R+ and GPT-4 Turbo following. Command R is substantially worse than R+ cross-lingually, unlike monolingually.
@@ 4.1 Impact of dataset
avg Okapi ShareGPT Complex (Ours)
@@ 4.5 Impact of instruction tuning
Table 6: Line-level pass rate of base vs instruction-tuned LLMs on monolingual generation for a subset of languages. Full results in Table A15, §A.8.
@@ 5 When does language confusion occur?
We study a sample of prompts to better understand where language confusion occurs. Intuitively, if a token in an undesired language is assigned sufficient probability, it may be sampled. We observe that language confusion typically occurs when the distribution over next tokens is flat and nucleus is large (see §A.9 for background).
@@ 6.1 Reducing temperature and nucleus size
We try to reduce the chance of having a wrong-language token sampled by manipulating temperature and nucleus size. Table 8 shows the results on monolingual WPR for Command R. Higher $T$ encourages language confusion: $T=1$ shows an average WPR of only 83.5%, and as low as 72.0% and 69.5% for Japanese and Chinese. Increasing $p$, resulting in a smaller nucleus, has a smaller effect.
@@ 6.2 Few-shot prompting
For cross-lingual generation, we use a similar format to the test sets. Figure A2 shows the template used to prompt the base model. For comparison, we also apply one-shot prompting to Command R where the example is presented as turns in a conversation. Table 9 shows the results.
@@ 6.3 Multilingual instruction tuning
We extend the English data with multilingual data Aryabumi et al. (2024). Due to the scarcity of multilingual data, our SFT data mixture is 90% English. For preference tuning, we use 50% multilingual data. Results are in Table 9.
@@ 7 Discussion
We have identified the issue of language confusion and demonstrated which model training and hyper-parameter choices impact it. However, there are several aspects that require further discussion.
Behavior of base models
English-centricity
Preference Tuning
Other factors
@@ 8 Related Work
Prior work has observed ‘source language hallucinations’ in zero-shot cross-lingual transfer Vu et al. (2022); Li and Murray (2023); Pfeiffer et al. (2023); Chirkova and Nikoulina (2024) when models are fine-tuned on English data and applied to generate text in another language. The problem of line-level language confusion is known in the machine translation field as ‘off-target translation’ Chen et al. (2023); Sennrich et al. (2024). It typically occurs on English-centric multilingual models, when used in a zero-shot manner (to translate between two languages unseen at training). For LLMs, there is no source language per se. A few studies Kew et al. (2023); Chen et al. (2024) have provided evidence of LLMs generating in an incorrect language on the response level. To the best of our knowledge, we are the first to show results on the sentence and word level and the first to systematically study language confusion in LLMs.
@@ 9 Conclusion
We have introduced the Language Confusion Benchmark to evaluate language confusion in LLMs. We have shown that some LLMs exhibit severe language confusion in many languages and even the strongest LLMs do not achieve perfect performance cross-lingually. We observed that base and English-centric instruct models are particularly susceptible to language confusion, which is exacerbated by complex prompts. Finally, we proposed measures to mitigate language confusion at inference (reducing temperature and nucleus size, few-shot prompting) and at training (multilingual instruction tuning). Our benchmark is efficient to evaluate and easy to extend and can help ensure that models achieve equal utility across languages.
@@ Acknowledgements
As this is to our knowledge, the first study on language confusion in LLMs, there are several natural research directions that we were not able to cover:
@@
Our metrics are applied to model outputs of at most 100 tokens. But what is a token depends on the model’s tokenizer. Models with more aggressive tokenization could be advantaged with regard to the binary metrics (LPR and WPR). Likewise, because we allow models to stop their generation early (by producing an end-of-sequence token), models which are less verbose could have an advantage.
@@ A.1 Extended Results
Table A3: Word-level pass rate (WPR) on monolingual and cross-lingual generation (of responses containing no English words). Llama models are instruction-tuned variants.

### v2

@@ 1 Introduction
Large language models (LLMs) are increasingly used in a variety of applications across the globe Kaddour et al. (2023). While early LLMs focused on English Joshi et al. (2020); Hu et al. (2020), recent models are more multilingual. Nevertheless, LLMs do not provide equal utility to non-English speakers due to higher latency, increased costs, and reduced performance Ahia et al. (2023); Asai et al. (2023); Held et al. (2023).
@@
Take an Arabic prompt as an example: an LLM may inappropriately respond fully in English (full-response confusion), produce some lines in the desired language and some in another language (line-level confusion), or sporadically insert single words or phrases in another language (word-level confusion). Figure 1 shows example errors. Even if such errors occur rarely, they cause a jarring user {CJK}UTF8gbsn经验 (experience).
We investigate language confusion on the line and word level in two practical settings: a) Monolingual generation, where a user queries the LLM in a given language, implicitly requesting an answer in the same language; and b) cross-lingual generation, where a user explicitly instructs a model to generate text in a different language.
@@ 2.2 Language Confusion Metrics
Off-the-shelf tools do not support word-level LID and LLMs only achieve 79–86 F1 detecting word-level code-switching Zhang et al. (2023), too low for use as automatic evaluators. Consequently, we take a two-pronged heuristic approach to detecting word-level language confusion focusing on settings where it achieves high precision. For non-Latin script languages, we observe that word-level language confusion in top LLMs mainly occurs with English. To avoid natural code-switching false positives, [^4] E.g., Japanese users may use English acronyms: {CJK}UTF8minAI に関する記事を書きます。(“Write an article about AI.”) we check for English words that do not typically occur in target language text. [^5] We source words from https://gist.github.com/WChargin/8927565, which is based on the Linux dictionary usually stored in /usr/share/dict/words. To reduce false positives, we exclude capitalized words from the dictionary (which are often proper nouns or acronyms). We evaluate word-level confusion in Arabic (ar), Hindi (hi), Japanese (ja), Korean (ko), Russian (ru), and Simplified Chinese (zh). For Latin script languages where word-level language confusion is rarer, [^6] We show an example of such word-level confusion in A1. we detect tokens where any character is outside of the Unicode range of the language’s script. We evaluate word-level confusion in German (de), English (en), Spanish (es), French (fr), Indonesian (id), Italian (it), Portuguese (pt), Turkish (tr), and Vietnamese (vi).
@@
A response is only correct when entirely in the correct language, as even one instance of language confusion can damage intelligibility and cause a jarring user experience. We calculate binary metrics to indicate whether a response contains any instance of a) a line in an incorrect language and b) an isolated English word for languages using $\neg$ Latin scripts and an out-of-Unicode-range character for Latin script languages. These main metrics are defined below.
@@
where $R$ is the set of all responses and $E_{L}$ the set of responses that contain line-level errors. [^7] For completeness, we show results for “line-level language accuracy” (the fraction of lines in the correct language across all responses of an LLM) in Tables A1 and A2 at §A.1.
Word-level pass rate (WPR): percentage of responses where all words are in the desired language. We exclude responses with line-level errors as most line-level errors would also be counted as word-level errors, making it difficult to disentangle the two error types. For languages that use a $\neg$ Latin script, we detect erroneous English words while for Latin script languages, we identify characters outside of the script’s Unicode range.
@@ 2.4 Data Filtering and Processing
As LID tools underperform on short sequences and non-standard text, we manually filter out: a) examples answerable with a single word/phrase; b) multiple-choice questions and prompts asking for lists; c) prompts requiring code generation, math equations, or data formats such as HTML. For datasets where completions are available, we filter out prompts with very short completions (less than 5 words).
@@
GPT-4o 99.7 98.1
Table 4: Average word-level pass rate (WPR) on non-Latin script languages. See Tables A3 and A4 for detailed WPR results on non-Latin and Latin script languages respectively.
@@ 3 Experiments
We evaluate the following LLMs covering various scales and model families: Llama 2 70B Instruct Touvron et al. (2023), Llama 3 and 3.1 70B Instruct Dubey et al. (2024), Command R (35B), [^12] https://cohere.com/blog/command-r Command R+ (104B parameters), [^13] https://cohere.com/blog/command-r-plus-microsoft-azure Command R Refresh (command-r-08-2024), Command R+ Refresh (command-r-plus-08-2024), Mixtral 8x7B Jiang et al. (2024), Mistral Large, [^14] https://mistral.ai/news/mistral-large/ GPT-3.5 Turbo (gpt-3.5-turbo-012524; Brown et al., 2020), GPT-4 Turbo (gpt-4-turbo-040924; Achiam et al., 2023), and GPT-4o (gpt-4o-2024-08-06). For Llama and Command models, we also evaluate base versions (see §4.5). We generate at most 100 tokens per prompt using nucleus sampling with $p=0.75$ and $T=0.3$.
LPR, WPR, and LCPR results are in Tables 3, 4, and A5, respectively. [^15] We find results show little variance across runs (see §A.9). We show language confusion examples for different models in Table A6.
@@
Command and GPT models perform well on average on the line level (LPR in $[98.6,99.3]$), but Llama 2 and 3 and Mistral models struggle to consistently generate text in the correct language (LPR in $[48.3,73.0]$). Llama 3.1 performs much better, however. Mistral Large is better on some European languages while Llama 2 and 3 exhibit language confusion even for high-resource languages such as German. Most models have WPR within the same range and perform better on Latin script vs non-Latin script languages; however, Mixtral 8x7B is considerably worse. GPT-4 Turbo is strongest on LCPR, with Command R+ Refresh and GPT-4o only slightly behind. Llama models’ preference for English responses leads to very low LCPR.
@@
In the challenging cross-lingual setting, the best models have LPRs in the low 90s. The Command R and R+ Refresh models outperform their original versions, with Command R Refresh in particular showing a large improvement over Command R, which performs poorly cross-lingually. OpenAI and Command models perform best; Command R+ Refresh achieves the best performance overall. Llama 2 and 3 models perform poorly: both scoring in the 30s due to a tendency to respond in English. Llama 3.1 shows improved performance but still tends to generate in English for some languages. Mistral Large is worse than Mixtral, even in European languages. On LCPR, Command R+ Refresh performs best, followed by GPT-4o.
@@ 4.1 Impact of dataset
Okapi ShareGPT Complex
avg (Ours)
@@ 4.5 Impact of instruction tuning
Table 6: Line-level pass rate (LPR) of base vs instruction-tuned LLMs on monolingual generation for a subset of languages. Full results in Table A17, §A.8.
@@ 5 When does language confusion occur?
We study a sample of prompts to better understand where language confusion occurs. Intuitively, if a token in an undesired language is assigned sufficient probability, it may be sampled. We observe that language confusion typically occurs when the distribution over next tokens is flat and the nucleus is large (see §A.10 for background).
@@ 6.1 Reducing temperature and nucleus size
[…]
@@
6.2 Beam search decoding
We explore the effect of beam search decoding on language confusion for Command R. Table 9 shows aggregate results for beam sizes 1 (greedy search) to 10, with full results in Section A.7. Increasing beam size helps moderately over greedy search on WPR, with little noticable effect on monolingual LPR. Increasing beam size consistently hurts cross-lingual LPR and is most pronounced for non-Indo-European languages, where average LPR drops from 73.9 with greedy search to 65.6 with beam size = 10.
Monoling. Crosslingual
WPR LPR WPR LPR
Overall $\neg$ IE IE
1 97.8 99.0 94.9 74.1 - 73.9 - 74.2 -
2 98.6 99.0 95.4 72.2 (-1.9) 71.1 (-2.8) 73.6 (-0.6)
3 98.6 98.7 97.1 71.5 (-2.5) 70.1 (-3.8) 73.4 (-0.9)
5 99.0 99.0 96.7 70.3 (-3.8) 68.3 (-5.6) 72.9 (-1.4)
10 99.0 98.5 96.7 68.4 (-5.7) 65.6 (-8.3) 72.1 (-2.1)
Table 9: Effect of beam search decoding on language confusion metrics for Command R. Beam sizes: 1-10.
Table 10 aggregates the best average score achieved for beam search and nucleus sampling from Tables 8, A13, A14, A15, A16. Beam search is always better than nucleus sampling for WPR and for cross-lingual LPR, and both methods seem equally effective for monolingual LPR, suggesting that beam search may be an effective decoding strategy for lessening language confusion (though at higher computational cost).
WPR LPR
Beam S. Nuc. Samp. Beam S. Nuc. Samp.
max min max min max min max min
Mono. 99.0 97.8 97.6 86.5 99.0 98.5 98.9 96.8
Cross. 97.1 94.9 94.6 80.1 74.1 68.4 68.2 64.0
Table 10: Beam Search vs. Nucleus Sampling. Shown: best and second-best overall average score.
@@ 6.3 Few-shot prompting
For cross-lingual generation, we use a similar format to the test sets. Figure A2 shows the template used to prompt the base model. For comparison, we also apply one-shot prompting to Command R where the example is presented as turns in a conversation. Table 11 shows the results. Tables A19 and A18 in Appendix show the detailed results per language for monolingual and cross-lingual language confusion respectively.
@@ 6.4 Multilingual instruction tuning
We extend the English data with multilingual data, most of which comes from machine-translated Dolly and ShareGPT Üstün et al. (2024). Due to the scarcity of multilingual data, our SFT data mixture is 90% English. For preference tuning, we use 50% multilingual data. Results are in Table 11.
@@ 7 Discussion
We discuss several aspects that raise important questions including the behavior of base models, the English centricity of models, the impact of preference tuning, and the importance of other factors in Appendix A.11.
@@ 8 Related Work
Prior work has observed ‘source language hallucinations’ in zero-shot cross-lingual transfer Vu et al. (2022); Li and Murray (2023); Pfeiffer et al. (2023); Chirkova and Nikoulina (2024) when models are fine-tuned on English data and applied to generate text in another language. The problem of language confusion is known in the machine translation field as ‘off-target translation’ Chen et al. (2023); Sennrich et al. (2024). It typically occurs on English-centric multilingual models, when used in a zero-shot manner (to translate between two languages unseen at training). For LLMs, there is no source language per se. A few studies Kew et al. (2023); Faisal and Anastasopoulos (2023); Chen et al. (2024) have provided evidence of LLMs generating in an incorrect language on the response level. Holtermann et al. (2024) analyze which languages are confused on the response level using mainly smaller LLMs. To the best of our knowledge, we are the first to show results on the line and word level and the first to systematically study language confusion in LLMs.
@@ 9 Conclusion
We have introduced the Language Confusion Benchmark. We have shown that some LLMs exhibit severe language confusion and even the strongest LLMs do not achieve perfect performance cross-lingually. We observed that base and English-centric instruct models are particularly susceptible to language confusion, which is exacerbated by complex prompts. Finally, we proposed measures to mitigate language confusion at inference and training. Our benchmark is efficient to evaluate and easy to extend and can help ensure that models achieve equal utility across languages.
@@ Acknowledgements
As this is to our knowledge, the first systematic study on language confusion in LLMs, there are several natural research directions that we were not able to cover:
@@
Our metrics are applied to model outputs of at most 100 tokens. But what is a token depends on the model’s tokenizer. Models with more aggressive tokenization could be advantaged with regard to the binary metrics (LPR and WPR). Likewise, because we allow models to stop their generation early (by producing an end-of-sequence token), models which are less verbose could have an advantage. Furthermore, because current LID tools do not support word-level language identification, our WPR metric is currently limited to assessing non-Latin script languages for unintended switching into English. As LID tools
@@ A.1 Extended Results
GPT-4o 99.7 100.0 100.0 100.0 99.0 99.0 100.0 98.1 99.2 97.0 97.5 99.1 98.0 97.6
Table A3: Word-level pass rate (WPR) on monolingual and cross-lingual generation in non-Latin script languages (of responses containing no English words). Llama models are instruction-tuned variants.
Monolingual Cross-lingual
Avg de en es fr id it pt tr vi Avg de es fr id it pt tr vi
Llama 3 70B-I 99.0 100.0 99.7 100.0 100.0 100.0 100.0 100.0 100.0 90.9 99.7 100.0 99.0 100.0 100.0 100.0 100.0 100.0 98.7
Llama 3.1 70B-I 99.6 100.0 99.7 100.0 100.0 98.9 100.0 100.0 99.0 99.0 99.1 100.0 99.2 99.6 99.1 100.0 99.6 98.0 97.5
Command R 99.2 100.0 99.7 100.0 100.0 100.0 100.0 100.0 96.9 96.0 97.9 98.9 100.0 100.0 96.4 100.0 98.5 97.4 92.2
Command R+ 99.6 100.0 100.0 100.0 100.0 99.0 100.0 100.0 100.0 97.0 98.6 100.0 100.0 100.0 98.6 99.6 100.0 97.9 93.1
Command R Refresh 99.9 100.0 100.0 99.7 100.0 100.0 100.0 99.5 100.0 100.0 99.9 100.0 100.0 100.0 99.5 100.0 100.0 100.0 100.0
Command R+ Refresh 99.9 100.0 99.9 100.0 100.0 100.0 100.0 100.0 100.0 99.0 99.7 100.0 100.0 100.0 99.1 100.0 100.0 99.0 99.6
GPT-3.5 Turbo 100.0 100.0 99.9 100.0 100.0 100.0 100.0 100.0 100.0 100.0 99.9 100.0 100.0 100.0 100.0 100.0 100.0 99.2 100.0
GPT-4o 99.8 100.0 99.7 100.0 100.0 98.9 100.0 99.5 100.0 100.0 99.7 100.0 100.0 100.0 99.1 100.0 100.0 99.6 99.1
Table A4: Word-level pass rate (WPR) on monolingual and cross-lingual generation in Latin script languages (of responses containing no characters that are outside the corresponding script’s Unicode range). Llama models are instruction-tuned variants.